How Many 2 Inch Square Blocks Can Fit Inside A 10inch By 20 Inch Box? how do u get the answer?
Found 2 solutions by mananth, ikleyn:
Answer by mananth(16949)

volume of 10 inch block = 10*10*10 cubic inches
=1000 cu.in
volume of 2" block = 2*2* cu. in
= 8 cu. in.
Number of 2" blocksa in 10" block = 1000 / 8= 250 blocks.
volume of 20" block = 20*20*20 = 8000 cu.in
number of 2" blocks = 8000/ 8 = 1000 blocks

Answer by ikleyn(53996)

.
How Many 2 Inch Square Blocks Can Fit Inside A 10inch By 20 Inch Box? how do u get the answer?
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~


        @mananth interprets this problem as 3D, although the problem does not give any reason for it.
        I will interpret and solve this problem as 2D for plane figures.


Place 5 blocks along the 10-inches dimension and 10 blocks along the 20-inches dimension.


In total, you will get 5 x 10 = 50 blocks.    ANSWER